OpenStudy (anonymous):

Manuallu calculate the compound interest on an investment of $12,000 at 7.5% interest, compounded semiannually, for 2 years.

OpenStudy (anonymous):

c.i=p(1+r/100)^n use this formula

OpenStudy (anonymous):

c.i-compound interest,p-principal,r=interest,n=time

OpenStudy (anonymous):

How would I do the problem

OpenStudy (anonymous):

p=12000,r=7.5,time is semiannual, so u take it as 4 12000(1+7.5/100)4=answer

OpenStudy (anonymous):

find out and tell me

OpenStudy (anonymous):

$51,600

OpenStudy (anonymous):

c.i=amount -money borrowed c.i=12000-answer=ur final answer

OpenStudy (anonymous):

ur calculation is wrong check it again

OpenStudy (anonymous):

its 16025.629 c.i=16025.629-12000=4025.629

OpenStudy (anonymous):

its power 4 not muliply it with 4

OpenStudy (anonymous):

answer is right?

OpenStudy (anonymous):

12000((1+7.5/100)^4)-1)=4025.629

OpenStudy (wolf1728):

First compute the semi-annual interest rate rate = (1+(.075/2))^2 annual rate = (1.0375)^2 annual rate = 1.07640625 annual rate = 7.640625% total = 12,000 * (1.07640625)^2 total = 13,903.81 And to double check that here is a calculator: http://1728.org/compint.htm

OpenStudy (wolf1728):

Oh and the interest earned is 13,903.81 minus 12,000.00 = 1,903.81

OpenStudy (anonymous):

no its wrong

OpenStudy (anonymous):

i need to know the meaning semiannual